OpenStudy (anonymous):

verify each identity: cos(alpha+beta)cos(alpha-beta)=cos^2beta-sin^2alpha

OpenStudy (anonymous):

\[\cos(\alpha+\beta)\cos(\alpha-\beta)=\cos^2\beta-\sin^2\alpha\]